OpenStudy (anonymous):

Lines A and B are represented by the equations given below: Line A: x + y = 5 Line B: x + y = 4 Which statement is true about the solution to the set of equations?

OpenStudy (anonymous):

A. It is (5, 4). B. It is (4, 5). C. There is no solution. D. There are infinitely many solutions.

OpenStudy (anonymous):

Can I please get an explanation on how to solve this as well?

OpenStudy (skullpatrol):

What do you think the answer is and why? :-)

OpenStudy (anonymous):

I'm not quite sure, my teacher has never gone over this that is why I'm here, do you think you could give an example and how to solve it? If I at least know how to solve this one I can post what i think my answer is.

OpenStudy (skullpatrol):

Replace x + y in the first equation by x + y from the second equation and tell me what you get.

OpenStudy (anonymous):

I think C, if not then D

OpenStudy (anonymous):

I don't see how it could be a or b seeing as It is two separate lines

OpenStudy (skullpatrol):

Does 5 equal 4?

OpenStudy (anonymous):

Nope

OpenStudy (skullpatrol):

x + y = 5 = x + y = 4

OpenStudy (anonymous):

A then?

OpenStudy (skullpatrol):

Nope.

OpenStudy (anonymous):

I still stand with C since they are both separate lines with no actual numbers in their equation other than the solution.

OpenStudy (skullpatrol):

You are correct by standing with C.

OpenStudy (skullpatrol):

They are parallel lines.

OpenStudy (anonymous):

Thank you for the help

OpenStudy (skullpatrol):

Thanks for asking :-)